Express the linear equation 3x-4y+20=0 in slope y intercept form.
Thank you,
LV

Answer by stanbon(75887)

Express the linear equation 3x-4y+20=0 in slope y intercept form.
---
3x - 4y = -20
---
4y = 3x + 20
---
y = (3/4)x + 5
==========================
Cheers,
Stan H.